OpenStudy (migitmack):

The surface area of a soup can represents the total area of the can's exterior and may be found by the formula S.A. = 2πrh + 2πr2. A soup company designing its own cans wants to determine the height of its cans with a given surface area and radius. Help the company by solving the surface area formula for h. PLZ HELP ME

OpenStudy (mimi_x3):

\[A = 2\pi rh + 2\pi r^2\]?

OpenStudy (migitmack):

where did u get that?

OpenStudy (mimi_x3):

from what you wrote lol is that right?

OpenStudy (migitmack):

no

OpenStudy (migitmack):

I jsut dont understand

OpenStudy (mimi_x3):

What do you understand?

OpenStudy (mimi_x3):

dont you*

OpenStudy (migitmack):

How to solve it

OpenStudy (mimi_x3):

Well, is the equation:\[\large A = 2\pi rh +2\pi r^{2}\]

OpenStudy (migitmack):

2*(3.14)*r*h+2(3.14)*r^2?

OpenStudy (migitmack):

If I do that I get the asnwer

OpenStudy (migitmack):

So i just put in the values?

OpenStudy (mimi_x3):

What? you are solving for \(h\)

OpenStudy (migitmack):

oh 1 sec

OpenStudy (migitmack):

i got the answer